# Consistency of Responses and Continuations Generated by Large Language Models on Social Media

### v2

@@ Conclusion
This study examined the semantic and emotional consistency of content generated by LLMs. Recognizing the emotional information embedded in text and the way LLMs handle these emotions is essential as they become more involved in human social contexts.
Our findings revealed that Gemma demonstrated a tendency to amplify negative emotions, particularly anger. Despite this bias, Gemma effectively perceived and transferred the original emotional tone of the text. On the other hand, the Llama model displayed stronger emotional retention across a broader spectrum, including anger, expectation, fear, optimism, and sadness, with fewer transitions to other emotions. For the semantic problem, both models show better performance for both continuation and response. LLMs can understand human input to some extent. These observations highlight distinct emotional dynamics between LLMs, offering insights for enhancing their design and application in emotion-sensitive contexts.
Acknowledgments
This study is partially supported by the 2023-2024 Annual Seed Research Funding of the Department of Science and Technology of Communication, University of Science and Technology of China (#ASRF2023).

### v3

@@ Conclusion
In this study, we evaluated how large language models handle emotional and semantic content in social-media contexts. By comparing four representative models—Gemma, Llama3, Claude, Llama3.3—on both continuation and response tasks, we uncovered distinct patterns in their emotional behaviors and contextual understanding.
Across both continuation and response tasks, LLMs tended to shift emotional content toward neutrality and generally reduced emotional intensity. This effect was more pronounced in response tasks, suggesting that LLMs favor emotionally moderate language when addressing existing posts. LLM-as-judge evaluations revealed that generated texts largely preserved thematic and logical coherence with source texts. While most models scored consistently high, variability in some (e.g., Claude) highlights sensitivity to prompt phrasing and model-specific robustness. The neutrality bias of LLMs could help defuse polarization and foster calmer dialogue around contentious topics. Conversely, the amplification of anger and fear poses both opportunities such as mobilization and risks like manipulation or echo-chamber reinforcement.
These findings reveal key emotional dynamics between LLMs, offering insights for improving their design and use in emotion-sensitive applications.